Write an equation for the parabola with the vertex (1,3) that passes through (-2,-15).
I have tried to figure out how to do y=A(x-h)^2, but cannot figure out how to apply it to the problem.
Sorry if topic is wrong, I can't click on the right one.

Answer by nerdybill(7384)

Write an equation for the parabola with the vertex (1,3) that passes through (-2,-15)
.
Vertex form of a parabola is
y = a(x-h)^2 + k
where
(h,k) is the vertex
.
Since the problem gives you the vertex we can plug it in:
y = a(x-1)^2 + 3
Now, we plug in the given point and solve for a:
-15 = a(-2-1)^2 + 3
-15 = a(-3)^2 + 3
-15 = a(9) + 3
-18 = a(9)
-2 = a
.
So, our equation is now:
y = -2(x-1)^2 + 3